# On the Analysis of Misspecified Variational Inequalities with Nonlinear Constraints

### v1

@@ 5 Numerical Experiment
$\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\sum_{j\in J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\left(\lambda_{k+1}-\lambda\right)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$
$\quad\geq-\left(\sum_{j\in J_{k}^{+}}\frac{\rho_{k}}{2}\bigl(f_{j}(x_{k+1},\theta^{*}\bigr)^{2}+\sum_{j\in J_{k}^{-}}\frac{\bigl(\lambda_{k}^{(j)}\bigr)^{2}}{2\rho_{k}}\right)$
$\quad\geq-\left(\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}+\rho_{k}L_{f,\theta}\|\theta_{k}-\theta^{*}\|\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1},\theta^{*})|\right),$ (14)
where the last inequality follows from (A). Therefore, we have
- $\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)+\sum_{j\in J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})+\left(\lambda_{k+1}-\lambda\right)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$
$\quad\leq\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}+\rho_{k}L_{f,\theta}\|\theta_{k}-\theta^{*}\|\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1},\theta^{*})|.$ (15)
Now using (A) and (A) we can show
$-\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)+\sum_{j\in J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})\leq\frac{1}{2\rho_{k}}\left(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\right)$
$\quad+L^{\Phi}_{\lambda\theta}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|+\rho_{k}L_{f,\theta}\|\theta_{k}-\theta^{*}\|\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1},\theta^{*})|.$ (16)
@@
Let $x\in X$ and $\lambda\geq 0$ be arbitrary vectors. From Algorithm 1 (line 4) we have:
@@
where in the last inequality, we use the convexity of $\Phi_{\rho_{k}}(\cdot,\lambda_{k},\theta^{*})$. Next, using the Cauchy-Schwarz inequality and Lipschitz continuity of $\nabla_{x}\Phi_{\rho_{k}}(x,\lambda,\cdot)$ with respect $x$ in $\theta$ with a constant $L^{\Phi}_{x\theta}$, we can show
@@
Moreover, using the three points equality, we can write:
@@
Using the fact that $\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})\leq\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})+\langle\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*}),x_{k+1}-x_{k}\rangle+\frac{L^{\Phi}}{2}\|x_{k+1}-x_{k}\|^{2}$, we can show
@@
$\qquad+\frac{L^{\Phi}}{2}\|x_{k+1}-x_{k}\|^{2}+\|x_{k+1}-x_{k}\|\Big(L_{F,x}\|x_{k}-x_{k-1}\|+L_{F,\theta}\|\theta_{k}-\theta_{k-1}\|\Big).$ (22)
Using (A) from Lemma (4.1), we can bound the left-hand side of (B) and obtain the following inequality
$\left(x_{k+1}-x\right)^{\top}F\left(x,\theta^{*}\right)+\sum_{j\in J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\Phi_{\rho_{k}}(x,\lambda_{k},\theta^{*})$
@@
$\qquad+L^{\Phi}_{\lambda\theta}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|+\frac{L^{\Phi}}{2}\|x_{k+1}-x_{k}\|^{2}+\rho_{k}L_{f,\theta}\|\theta_{k}-\theta^{*}\|\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1},\theta^{*})|.$ (23)
Note that, applying Young’s inequality to the term $\|x_{k+1}-x_{k}\|\Big(L_{F,x}\|x_{k}-x_{k-1}\|+L_{F,\theta}\|\theta_{k}-\theta_{k-1}\|\Big)$, we have
@@
Now let’s add and subtract $\frac{L_{F,x}}{2}\|x_{k+1}-x_{k}\|^{2}$ and $\frac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$ with the respective terms to obtain
@@
Putting all terms together in (B), we get
$\left(x_{k+1}-x\right)^{\top}F\left(x,\theta^{*}\right)+\sum_{j\in J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\Phi_{\rho_{k}}(x,\lambda_{k},\theta^{*})$
$\quad\leq\frac{1}{2\gamma_{k}}\Big(\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}\Big)+\frac{1}{2\rho_{k}}\Big(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\Big)$
$\qquad+\left(\frac{L^{\Phi}+2L_{F,x}+L_{F,\theta}}{2}-\frac{1}{2\gamma_{k}}\right)\|x_{k+1}-x_{k}\|^{2}+\frac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}+\langle r_{k+1},x_{k+1}-x\rangle$
$\qquad-\langle r_{k},x_{k}-x\rangle+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x\|\Big)$
@@
$\qquad+L_{\lambda\theta}^{\Phi}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|+\rho_{k}L_{f,\theta}\|\theta_{k}-\theta^{*}\|\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1},\theta^{*})|.$ (24)
Appendix C Proof of Lemma 4.3
@@
From (B), setting $(x,\lambda)=(x^{*},\lambda^{*})$ gives, for all $k\geq 0$
$0\leq\frac{1}{2\gamma}\Big(\|x_{k}-x^{*}\|^{2}-\|x_{k+1}-x^{*}\|^{2}\Big)+\left(\frac{L^{\Phi}+2L_{F,x}+L_{F,\theta}}{2}-\frac{1}{2\gamma}\right)\|x_{k+1}-x_{k}\|^{2}$
$\quad+\frac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}+\langle r_{k+1},x_{k+1}-x^{*}\rangle-\langle r_{k},x_{k}-x^{*}\rangle+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x^{*}\|$
$\quad+L_{\lambda\theta}^{\Phi}\|\lambda_{k+1}-\lambda^{*}\|\|\theta_{k}-\theta^{*}\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x^{*}\|\Big)$
$\quad+\frac{L_{F,x}}{2}\left(\|x_{k}-x_{k-1}\|^{2}-\|x_{k+1}-x_{k}\|^{2}\right)+\frac{L_{F,\theta}}{2}\left(\|\theta_{k}-\theta_{k-1}\|^{2}-\|\theta_{k+1}-\theta_{k}\|^{2}\right)$
$\quad+\frac{1}{2\rho}\Big(\|\lambda_{k}-\lambda^{*}\|^{2}-\|\lambda_{k+1}-\lambda^{*}\|^{2}\Big)+\rho L_{f,\theta}\|\theta_{k}-\theta^{*}\|\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1},\theta^{*})|.$ (25)
By the choice of $\gamma\in\Big(0,\ (L^{\Phi}+2L_{F,x}+L_{F,\theta})^{-1}\Big)$, the coefficient of $\|x_{k+1}-x_{k}\|^{2}$ in (C) is non-positive, and discarding that term yields
$0\leq\frac{1}{2\gamma}\Big(\|x_{k}-x^{*}\|^{2}-\|x_{k+1}-x^{*}\|^{2}\Big)+\frac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}+\langle r_{k+1},x_{k+1}-x^{*}\rangle-\langle r_{k},x_{k}-x^{*}\rangle$
$\quad+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x^{*}\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x^{*}\|\Big)$
$\quad+L_{\lambda\theta}^{\Phi}\|\lambda_{k+1}-\lambda^{*}\|\|\theta_{k}-\theta^{*}\|+\frac{L_{F,x}}{2}\left(\|x_{k}-x_{k-1}\|^{2}-\|x_{k+1}-x_{k}\|^{2}\right)$
$\quad+\frac{L_{F,\theta}}{2}\left(\|\theta_{k}-\theta_{k-1}\|^{2}-\|\theta_{k+1}-\theta_{k}\|^{2}\right)+\frac{1}{2\rho}\Big(\|\lambda_{k}-\lambda^{*}\|^{2}-\|\lambda_{k+1}-\lambda^{*}\|^{2}\Big)$
$\quad+\rho L_{f,\theta}\|\theta_{k}-\theta^{*}\|\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1},\theta^{*})|.$ (26)
By applying Young’s inequality to the cross term in (C) for any $\alpha_{k}>0$, we have

### v3

@@ 5 Numerical Experiment
$\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)-\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\left(\lambda_{k+1}-\lambda\right)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$
$\quad\geq-\left(\sum_{j\in J_{k}^{+}}\frac{\rho_{k}}{2}\bigl(f_{j}(x_{k+1},\theta^{*}\bigr)^{2}+\sum_{j\in J_{k}^{-}}\frac{\bigl(\lambda_{k}^{(j)}\bigr)^{2}}{2\rho_{k}}\right)+E_{k}$
$\quad\geq-\left(\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}\right)+E_{k},$ (14)
where the last inequality follows from (11). Therefore, we have
- $\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})+\left(\lambda_{k+1}-\lambda\right)^{\top}\nabla_{\lambda}\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})$
$\quad\leq\left(\frac{1}{2\rho_{k}}\|\lambda_{k+1}-\lambda_{k}\|^{2}+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}\right)-E_{k}.$ (15)
Recall that $E_{k}=\sum_{j\in J_{k}^{+}}\left(\frac{\rho_{k}}{2}\Bigl(f_{j}(x_{k+1},\theta_{k})^{2}-f_{j}(x_{k+1},\theta^{*})^{2}\Bigr)+\lambda_{k}^{(j)}\Bigl(f_{j}(x_{k+1},\theta_{k})-f_{j}(x_{k+1},\theta^{*})\Bigr)\right)+\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})-\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k}).$ Using the triangle inequality and $a^{2}-b^{2}=(a-b)(a+b)$, we obtain
$|E_{k}|$ $\leq\sum_{j\in J_{k}^{+}}\frac{\rho_{k}}{2}|f_{j}(x_{k+1},\theta_{k})-f_{j}(x_{k+1},\theta^{*})||f_{j}(x_{k+1},\theta_{k})+f_{j}(x_{k+1},\theta^{*})|$
$\qquad+\sum_{j\in J_{k}^{+}}|\lambda_{k}^{(j)}||f_{j}(x_{k+1},\theta_{k})-f_{j}(x_{k+1},\theta^{*})|+\left|\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})-\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta_{k})\right|.$ (16)
Recall that $\Phi_{\rho}(x,\lambda,\theta)=\sum_{j=1}^{J}\phi_{\rho}(f_{j}(x,\theta),\lambda_{j})$, moreover, one can observe that for any $u_{j},u_{j}^{\prime}\in\mathbb{R}$, $|\phi_{\rho}(u_{j},v_{j})-\phi_{\rho}(u_{j}^{\prime},v_{j})|\leq(|v_{j}|+\rho D_{f})|u_{j}-u_{j}^{\prime}|$. Therefore, by Lipschitz continuity of $f_{j}(x,\theta)$ in $\theta$ we conclude that for any $\theta,\theta^{\prime}\in\Theta$, $|\phi_{\rho}(f_{j}(x,\theta),\lambda^{(j)})-\phi_{\rho}(f_{j}(x,\theta^{\prime}),\lambda^{(j)})|\leq(|\lambda^{(j)}|+\rho D_{f})L_{f,\theta}\|\theta-\theta^{\prime}\|$ which summing over $j$ implies that
$|\Phi_{\rho}(x,\lambda,\theta)-\Phi_{\rho}(x,\lambda,\theta^{\prime})|\leq L_{f,\theta}(J\rho D_{f}+\sqrt{J}\|\lambda\|)\|\theta-\theta^{\prime}\|.$ (17)
Using the above inequality within (A) and by Lipschitz continuity of $f_{j}(x,\theta)$ in $\theta$ and the bound $|f_{j}(x,\theta)|\leq D_{f}$, we conclude that
$|E_{k}|\leq 2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$ (18)
Now using (A) and (18) in (A), we can show
$-\Phi_{\rho_{k}}\bigl(x_{k+1},\lambda_{k},\theta^{*}\bigr)+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})$ $\leq\frac{1}{2\rho_{k}}\left(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\right)+L^{\Phi}_{\lambda\theta}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|$
$\quad+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}+2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$ (19)
@@
Proof Recall from Definition 2.1 and (4), for any $x\in X$, we have
$\nabla_{x}\Phi_{\rho}(x,\lambda,\theta)=\sum_{j=1}^{J}[\rho f_{j}(x,\theta)+\lambda^{(j)}]_{+}\,\nabla_{x}f_{j}(x,\theta)=\nabla_{x}f(x,\theta)^{\top}[\rho f(x,\theta)+\lambda]_{+}.$
Fix $\rho>0$, $\lambda\in\mathbb{R}^{J}$, and $\theta\in\Theta$. Then for any $x,y\in X$, we have
$\nabla_{x}\Phi_{\rho}(x,\lambda,\theta)-\nabla_{x}\Phi_{\rho}(y,\lambda,\theta)$ $=\nabla_{x}f(x,\theta)^{\top}[\rho f(x,\theta)+\lambda]_{+}-\nabla_{x}f(y,\theta)^{\top}[\rho f(y,\theta)+\lambda]_{+}$
$=\big(\nabla_{x}f(x,\theta)-\nabla_{x}f(y,\theta)\big)^{\top}[\rho f(x,\theta)+\lambda]_{+}$
$\quad+\nabla_{x}f(y,\theta)^{\top}\Big([\rho f(x,\theta)+\lambda]_{+}-[\rho f(y,\theta)+\lambda]_{+}\Big),$
where we added and subtracted $\nabla_{x}f(y,\theta)^{\top}[\rho f(x,\theta)+\lambda]_{+}$. Taking norms and applying the triangle inequality yields
$\|\nabla_{x}\Phi_{\rho}(x,\lambda,\theta)-\nabla_{x}\Phi_{\rho}(y,\lambda,\theta)\|$ $\leq\|\nabla_{x}f(x,\theta)-\nabla_{x}f(y,\theta)\|\,\|[\rho f(x,\theta)+\lambda]_{+}\|$
$\quad+\|\nabla_{x}f(y,\theta)\|\,\|[\rho f(x,\theta)+\lambda]_{+}-[\rho f(y,\theta)+\lambda]_{+}\|.$ (20)
Since $\cdot$$_{+}$ is the projection onto $\mathbb{R}^{J}_{+}$, it is non-expansive. Therefore,
$\|f(x,\theta)-f(y,\theta)\|=\Big(\sum_{j=1}^{J}|f_{j}(x,\theta)-f_{j}(y,\theta)|^{2}\Big)^{1/2}\leq\sqrt{J}\,L_{f,x}\|x-y\|.$
Consequently, it follows that
$\|[\rho f(x,\theta)+\lambda]_{+}-[\rho f(y,\theta)+\lambda]_{+}\|\leq\rho\sqrt{J}\,L_{f,x}\|x-y\|.$
Similarly, by Assumption 2.1(iii), each $\nabla_{x}f_{j}(\cdot,\theta)$ is Lipschitz continuous in $x$ with constant $L_{\nabla f,x}$. Therefore, the Jacobian matrix satisfies
$\|\nabla_{x}f(x,\theta)-\nabla_{x}f(y,\theta)\|\leq\sqrt{J}\,L_{\nabla f,x}\|x-y\|.$
Substituting these bounds into (B), we obtain
$\|\nabla_{x}\Phi_{\rho}(x,\lambda,\theta)-\nabla_{x}\Phi_{\rho}(y,\lambda,\theta)\|$ $\leq\sqrt{J}\,L_{\nabla f,x}\|x-y\|\,\|[\rho f(x,\theta)+\lambda]_{+}\|$
$\quad+\rho\sqrt{J}\,L_{f,x}\|x-y\|\,\|\nabla_{x}f(y,\theta)\|.$
Now, since $X\times\Theta$ is compact and the mappings $(z,\theta)\mapsto f(z,\theta)$ and $(z,\theta)\mapsto\nabla_{x}f(z,\theta)$ are continuous, the quantities $D_{f}\triangleq\sup_{(z,\vartheta)\in X\times\Theta}\|[f(z,\vartheta)]_{+}\|<\infty,\quad M_{\nabla f}\triangleq\sup_{(z,\vartheta)\in X\times\Theta}\|\nabla_{x}f(z,\vartheta)\|<\infty,\$are well defined and finite. Moreover, $\|[\rho f(x,\theta)+\lambda]_{+}\|\leq\rho\|[f(x,\theta)]_{+}\|+\|\lambda\|\leq\rho D_{f}+\|\lambda\|,$ and $\|\nabla_{x}f(y,\theta)\|\leq M_{\nabla f}$. Hence, we get
$\|\nabla_{x}\Phi_{\rho}(x,\lambda,\theta)-\nabla_{x}\Phi_{\rho}(y,\lambda,\theta)\|$ $\leq\sqrt{J}\,L_{\nabla f,x}(\rho D_{f}+\|\lambda\|)\|x-y\|+\rho\sqrt{J}\,L_{f,x}M_{\nabla f}\|x-y\|.$
Rearranging terms gives
$\|\nabla_{x}\Phi_{\rho}(x,\lambda,\theta)-\nabla_{x}\Phi_{\rho}(y,\lambda,\theta)\|\leq\Big(\rho C_{1}+\|\lambda\|C_{2}\Big)\|x-y\|,$
where $C_{1}\triangleq\sqrt{J}\big(L_{\nabla f,x}D_{f}+L_{f,x}M_{\nabla f}\big),\quad C_{2}\triangleq\sqrt{J}\,L_{\nabla f,x}.$ Thus, $\nabla_{x}\Phi_{\rho}(\cdot,\lambda,\theta)$ is Lipschitz continuous on $X$. ∎
Appendix C Proof of Lemma 4.3
Proof.
Proof Let $x\in X$ and $\lambda\in\mathbb{R}^{J}_{+}$ be arbitrary vectors. From Algorithm 1 (line 5) we have:
@@
where in the last inequality, we use the convexity of $\Phi_{\rho_{k}}(\cdot,\lambda_{k},\theta^{*})$. Next, using the Cauchy-Schwarz inequality and Lipschitz continuity of $\nabla_{x}\Phi_{\rho_{k}}(x,\lambda,\cdot)$ with respect $\theta$ with a constant $L^{\Phi}_{x\theta}$, we can show
@@
Moreover, using the three-point equality, we can write:
@@
Next, using the Lemma 4.2, we have $\Phi_{\rho_{k}}(x_{k+1},\lambda_{k},\theta^{*})\leq\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*})+\langle\nabla_{x}\Phi_{\rho_{k}}(x_{k},\lambda_{k},\theta^{*}),x_{k+1}-x_{k}\rangle+\frac{\rho_{k}C_{1}+\|\lambda_{k}\|C_{2}}{2}\|x_{k+1}-x_{k}\|^{2}$. Substituting this estimate into (C) yields
@@
$\qquad+\frac{\rho_{k}C_{1}+\|\lambda_{k}\|C_{2}}{2}\|x_{k+1}-x_{k}\|^{2}+\|x_{k+1}-x_{k}\|\Big(L_{F,x}\|x_{k}-x_{k-1}\|+L_{F,\theta}\|\theta_{k}-\theta_{k-1}\|\Big).$ (26)
To treat the term involving $\|\lambda_{k}\|$, we first add and subtract $\lambda^{*}$ inside the norm and use the triangle inequality to obtain
$\frac{\rho_{k}C_{1}+\|\lambda_{k}\|C_{2}}{2}\|x_{k+1}-x_{k}\|^{2}$ $=\left(\frac{\rho_{k}C_{1}+\|\lambda^{*}+(\lambda_{k}-\lambda^{*})\|C_{2}}{2}\right)\|x_{k+1}-x_{k}\|^{2}$
$\leq\frac{\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}}{2}\|x_{k+1}-x_{k}\|^{2}+\frac{C_{2}}{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}.$ (27)
Using (A) from Lemma (4.1) and substituting (C) into (C), we obtain
$\left(x_{k+1}-x\right)^{\top}F\left(x,\theta^{*}\right)+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\Phi_{\rho_{k}}(x,\lambda_{k},\theta^{*})$
$\quad\leq\frac{1}{2\gamma_{k}}\Big(\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}-\|x_{k+1}-x_{k}\|^{2}\Big)+\langle r_{k+1},x_{k+1}-x\rangle-\langle r_{k},x_{k}-x\rangle+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}$
$\qquad+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x\|\Big)+L^{\Phi}_{\lambda\theta}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|$
@@
$\qquad+\frac{\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}}{2}\|x_{k+1}-x_{k}\|^{2}+\frac{C_{2}}{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}+2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$ (28)
Note that applying Young’s inequality to the term $\|x_{k+1}-x_{k}\|\Big(L_{F,x}\|x_{k}-x_{k-1}\|+L_{F,\theta}\|\theta_{k}-\theta_{k-1}\|\Big)$, we have
@@
Now let us add and subtract $\frac{L_{F,x}}{2}\|x_{k+1}-x_{k}\|^{2}$ and $\frac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$ with the respective terms to obtain
@@
Putting all terms together in (C), we get
$\left(x_{k+1}-x\right)^{\top}F\left(x,\theta^{*}\right)+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\Phi_{\rho_{k}}(x,\lambda_{k},\theta^{*})$
$\quad\leq\frac{1}{2\gamma_{k}}\Big(\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}\Big)+\frac{1}{2\rho_{k}}\Big(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\Big)+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}$
$\qquad+\left(\frac{\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}}{2}-\frac{1}{2\gamma_{k}}\right)\|x_{k+1}-x_{k}\|^{2}+\frac{C_{2}}{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}+\frac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
$\qquad+\langle r_{k+1},x_{k+1}-x\rangle-\langle r_{k},x_{k}-x\rangle+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x\|\Big)$
@@
$\qquad+L_{\lambda\theta}^{\Phi}\|\lambda_{k+1}-\lambda\|\|\theta_{k}-\theta^{*}\|+2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$ (29)
Appendix D Proof of Lemma 4.4
@@
Proof From (C), setting $(x,\lambda)=(x^{*},\lambda^{*})$ gives, for all $k\geq 0$
$0\leq\frac{1}{2\gamma_{k}}\Big(\|x_{k}-x^{*}\|^{2}-\|x_{k+1}-x^{*}\|^{2}\Big)+\frac{1}{2\rho_{k}}\Big(\|\lambda_{k}-\lambda^{*}\|^{2}-\|\lambda_{k+1}-\lambda^{*}\|^{2}\Big)+\frac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
$\qquad+\left(\frac{\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}}{2}-\frac{1}{2\gamma_{k}}\right)\|x_{k+1}-x_{k}\|^{2}+\frac{C_{2}}{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}-\langle r_{k},x_{k}-x^{*}\rangle$
$\qquad+\langle r_{k+1},x_{k+1}-x^{*}\rangle+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x^{*}\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\Big(\|x_{k+1}-x_{k}\|+\|x_{k}-x^{*}\|\Big)$
$\qquad+L_{\lambda\theta}^{\Phi}\|\lambda_{k+1}-\lambda^{*}\|\|\theta_{k}-\theta^{*}\|+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}+2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$ (30)
To bound the cross term $L_{\lambda\theta}^{\Phi}\|\lambda_{k+1}-\lambda^{*}\|\|\theta_{k}-\theta^{*}\|$, we apply Young’s inequality for any $\alpha_{k}>0$